# bibfile.bib

### v1

@@ 3.2 Property (FH) and soluble normal subgroups
The main tool to prove Theorem 3.4 will be a form of Levi decomposition for parabolic subgroups of $H=\text{Isom}(X)$, where a parabolic subgroup will be expressed as a semidirect product of a soluble normal subgroup and a Levi factor of lower rank. In this subsection, we prove a lemma that will allow us to restrict the action to the Levi factor, and thus to reduce the rank of the target space.
@@
Let $N^{\prime}:=[N,N]$ be the commutator subgroup of $N$, which has solubility length $n-1$, let $A=N/N^{\prime}$ be the abelianization of $A$, and let $ab:N\rightarrow A$ be the natural projection (with kernel $N^{\prime}$). We recall that $Q$ acts on $N$ by the map $\phi:Q\rightarrow\text{Aut}(N)$ that appears in the semidirect product and that associates to every $q\in Q$ the conjugation by $q$. Since the commutator subgroup $N^{\prime}$ is characteristic, it is preserved by the action of $Q$, which means that the action descends to the quotient $A$. We call $\phi_{ab}$ this induced action. By hypothesis, $A$ is isomorphic (as a topological group) to the additive group of a Hilbert space, and the action $\phi_{ab}$ of $Q$ preserves a scalar product in this Hilbert space.
@@
Using Property (FH), we find that the action of $\Gamma$ on $A$ fixes a point $a_{0}$, which means that if we conjugate $f$ by an element $n_{0}\in N$ with $ab(n_{0})=a_{0}$ we get an action that fixes the neutral element of $A$, which means that a conjugate of the homomorphism $f$ has image contained in $N^{\prime}\rtimes Q$. Now, since $N^{\prime}$ has solubility length $n-1$, by inductive hypothesis this conjugate has image contained in a conjugate of $Q$, so in the end we find that $f$ also has image contained in a conjugate of $Q$.
@@ 3.3 The induction in the finite-dimensional case
As the group $\Gamma$ has Property (F), we know that every action of $\Gamma$ on $X$ has to fix a point either in $X$ or in $\partial X$. The first case is the thesis of our theorem, so from now on we will assume that we are in the second case, i. e. $\Gamma$ fixes $x\in\partial X$. This means that there is a homomorphism $f:\Gamma\rightarrow H_{x}$, where $H_{x}<H=\text{Isom}(X)$ is the parabolic subgroup defined by $H_{x}=\{h\in H:hx=x\}$.
In this subsection, we focus on the case where $X$ is finite-dimensional. We will use the classical theory of Lie groups and symmetric spaces of non-compact type, for which a good reference is the book by Eberlein [eberlein]. We use results from (2.17) of that book, which we report here using our notation.
@@
Eberlein then defines $\mathfrak{n}_{x}$ as the Lie algebra of $N_{x}$ and shows that it is nilpotent (2.17.14), which means that the group $N_{x}$ is nilpotent in the algebraic sense ([hilneeb], Theorem 11.2.5). Moreover, as shown in the proof of the Iwasawa decomposition ([hilneeb], Theorem 13.3.8), $N_{x}$ is simply connected.
Now, if we have the homomorphism $f:\Gamma\rightarrow H_{x}$, we can compose it with the projection $\pi:H_{x}\rightarrow Z_{x}$ and get a homomorphism $\overline{f}:\Gamma\rightarrow Z_{x}$. In the proof of ([eberlein], 2.17.5) it is shown that if $\gamma_{px}$ is a geodesic that passes through $p$ and points to $x$, the set $F(\gamma_{px})$ defined as the union of all geodesics parallel to $\gamma_{px}$ is preserved by the action of $Z_{x}$. Therefore, we have an action of $\Gamma$ on $F(\gamma_{px})$. It is shown in [eberlein] (2.11.4) that $F(\gamma_{px})$ is a complete totally geodesic submanifold of $X$ and
@@
As $\Gamma$ acts on $F(\gamma_{px})$, the result of De Rham [derham] implies that the action splits as the product of an action on $\mathbb{R}^{r}$ and an action on $F_{S}(\gamma_{px})$. Due to Property (FH), the action on $\mathbb{R}^{r}$ fixes a point; hence there exists an embedded copy of $F_{S}(\gamma_{px})$ that is preserved by $\Gamma$ and on which $\Gamma$ acts.
As $\Gamma$ has Property (F) it must fix a point either in $F_{S}(\gamma_{px})$ or in its boundary. Since $F(\gamma_{px})$ is embedded in $X$ and since $r\geq 1$, the rank of $F_{S}(\gamma_{px})$ must be strictly lower than the rank of $X$, so we can use our inductive hypothesis and find that $\Gamma$ has to fix a point inside $F_{S}(\gamma_{px})$.
@@
To verify the hypotheses of Lemma 3.5, we need to see that each Abelian quotient that appears in the derived series of the soluble group $N_{x}$ is isomorphic to the additive group of a Hilbert space where the action of $Q$ preserves a scalar product. We know from [hilneeb] (Proposition 11.2.4) that the groups in the derived series of a connected Lie group $\mathcal{G}$ are integral subgroups (that is, they are generated by a subalgebra of the Lie algebra of $\mathcal{G}$), and that integral subgroups $\mathcal{H}$ of a simply connected soluble Lie group $\mathcal{G}$ are closed and simply connected, with $\mathcal{G}/\mathcal{H}$ isomorphic to $\mathbb{R}^{\dim{\mathcal{G}/\mathcal{H}}}$ ([hilneeb], Proposition 11.2.15). We can apply this to the Lie group $\mathcal{G}=N_{x}$, finding that the groups in the derived series are closed and simply connected, and each quotient is isomorphic to $\mathbb{R}^{n}$ for some $n$. We also know that the groups in the derived series are characteristic subgroups of $N_{x}$ (meaning that they are preserved by any automorphism of $N_{x}$), and hence the action of $Q$ given by the semidirect product $N_{x}\rtimes Q$ descends to an action on each quotient. Furthermore, since $Q$ is compact, each quotient has a scalar product that is preserved by this action of $Q$: it suffices to start from any scalar product and integrate it over $Q$ to obtain a $Q$-invariant one. This completes the proof.
@@ 3.4 The infinite-dimensional case
We are left with showing the inductive step for the infinite-dimensional case, where we have $X=X_{p}(\mathbb{K})$ for some $p\geq 1$ and $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$. Since $\Gamma$ acts by isometries on $X$, it has a homomorphism to the group $H=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, where $\mathrm{PO}_{\mathbb{R}}(p,\infty)=\mathrm{PO}(p,\infty)$, $\mathrm{PO}_{\mathbb{C}}(p,\infty)=\mathrm{PU}(p,\infty)$, and $\mathrm{PO}_{\mathbb{H}}(p,\infty)=\mathrm{PSp}(p,\infty)$. In fact, it is shown in [duc2023] (Theorem 3.3) that $\mathrm{Isom}(X_{p}(\mathbb{K}))=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, except in the case $\mathbb{K}=\mathbb{C}$ where $\mathrm{PU}(p,\infty)$ is a subgroup of index two in $\mathrm{Isom}(X_{p}(\mathbb{C}))$. In that case, we have a subgroup $\Gamma^{\prime}<\Gamma$ of index $\leq 2$ with a homomorphism to $\mathrm{PU}(p,\infty)$, and if we show that $\Gamma^{\prime}$ fixes a point in $X$, then this implies that $\Gamma$ has a finite orbit and hence fixes a point.
@@
Each point at infinity of $X_{p}(\mathbb{K})$ is associated with a flag of isotropic subspaces of the vector space $V$: this was shown in [duc2013] (Proposition 6.1) in the real case, but is true in all cases, as observed in [duc2023] (as a consequence of Remark 3.1 in [duc2023]).
We consider the flag associated with the point at infinity $x$ (which is fixed by the action of $\Gamma$). Let $q\leq p$ be the dimension of the maximal subspace in the flag. We may choose the basis in such a way that this maximal subspace is exactly $\text{Span}\{e_{1},\ldots,e_{q}\}$. Then we can change the order of the elements $(e_{q+1},e_{q+2},\ldots)$ of the basis in such a way that the sesquilinear form is now expressed by the matrix
@@
Now we consider the matrix associated with a transformation in $H_{x}$ with respect to this new basis, and find some restrictions on its structure. We see it as a block matrix similar to the one expressing the sesquilinear form, i. e. with two blocks of dimension $q$ and one block of infinite dimension, both horizontally and vertically. This means that we are splitting the vector space $V$ into a direct sum $V_{1}\oplus V_{2}\oplus V_{3}$ where $V_{1}$ and $V_{2}$ have dimension $q$ and $V_{3}$ has infinite dimension, and each block of the matrix corresponds to a continuous linear map from $V_{i}$ to $V_{j}$ for some $i,j\in\{1,2,3\}$.
@@
In order to preserve the sesquilinear form we must have
@@
The map sending a matrix as above to the block $B$ determines a homomorphism $N\rightarrow A$ where $A$ is the additive group of $q\times\infty$ matrices (i. e. linear maps from the $q$-dimensional vector space $V_{2}$ to the infinite-dimensional Hilbert space $V_{3}$), which is Abelian and isomorphic with the direct sum of $q$ Hilbert spaces of countable dimension, which is isomorphic with a single Hilbert space of countable dimension. The kernel of the map is the set of matrices in $N$ with $B=0$, which is isomorphic with the additive group of $q\times q$ matrices $Y$ that satisfy $Y+Y^{T}=0$, i. e. the additive group of skew-symmetric matrices, isomorphic with $\mathbb{K}^{q(q-1)/2}$.
@@
What we need to show for our theorem is that any continuous homomorphism $f:\Gamma\rightarrow H_{x}$ has image contained in the stabilizer of a point in the symmetric space $X$. Consider the homomorphisms $f_{S}:\Gamma\rightarrow\mathrm{P}S$ and $f_{O}:\Gamma\rightarrow\mathrm{PO}_{\mathbb{K}}(p-q,\infty)$ defined by $f_{S}=\pi_{S}\circ f$ and $f_{O}=\pi_{O}\circ f$, where $\pi_{S}$ and $\pi_{O}$ are the natural projections of $H_{x}$ onto $\mathrm{P}S$ and $\mathrm{PO}_{\mathbb{K}}(p-q,\infty)$ respectively.
The homomorphism $f_{S}$ has image in $\mathrm{P}S<\mathrm{PGL}_{q}(\mathbb{K})$. As $\mathrm{PGL}_{q}(\mathbb{K})$ is a Lie group of rank $q-1<p$, we can use the inductive hypothesis of Theorem 3.4 and see that the image of $f_{S}$ fixes a point in the associated symmetric space. Up to conjugating the homomorphism (which is the same as changing the basis of $V_{1}$ and $V_{2}$ that we consider in our matrix decomposition), we may assume that the image of $f_{S}$ is contained in $\mathrm{PO}_{\mathbb{K}}(q)$.
@@ 3.5 Conclusion
We are now ready to prove Theorem 1.4.
Let $G=\mathrm{SL}_{3}(\mathbb{F})$, and let $\Gamma$ be a discrete, torsion-free, cocompact subgroup of $G$, which exists by ([borelharder], Theorem A, Theorem 3.3) and ([garland], Theorem 2.7). By Theorem 2.1, $\Gamma$ has Property (F). Since finite-index subgroups of $\Gamma$ are again discrete, torsion-free, and cocompact in $G$, they also have Property (F).
The group $G$ has Kazhdan’s Property (T) ([kazhdanT], 1.4), and since this property is inherited by lattices ([kazhdanT], Theorem 1.7.1), the subgroup $\Gamma$ and all its finite-index subgroups also have Kazhdan’s Property (T). By the Delorme-Guichardet Theorem ([kazhdanT], Theorem 2.12.4), this means that these groups have Property (FH).
By Theorem 3.4, any continuous action by isometries of $\Gamma$ on a space $X\in\mathcal{X}$ has a fixed point in $X$. As explained in Remark 1.6, this implies that any continuous action by isometries of $G$ on $X$ has a fixed point, proving Theorem 1.4.
@@ Appendix A Appendix
In this appendix we prove that the spaces $X_{p}(\mathbb{K})$, as defined at the beginning of Section 3.1, are irreducible metric spaces, meaning that they are not isometric to the product of two metric spaces not reduced to a point.
Set $X=X_{p}(\mathbb{K})$. We suppose by contradiction that $X=X_{1}\times X_{2}$ for some metric spaces $X_{1},X_{2}$ not reduced to a point. First, we note that $X_{1},X_{2}$ must be geodesic spaces, since $X$ is a geodesic space and geodesics in a product project to geodesics in factors ([bh], I.5.3). Moreover, geodesics in $X_{1},X_{2}$ must extend indefinitely, since geodesics in $X$ extend indefinitely. This implies that the boundary at infinity of $X_{1}$ and $X_{2}$ is non-empty. We also note that $X_{1},X_{2}$ must be CAT(0) spaces, as they are convex subsets of the CAT(0) space $X$, and that they must be complete, as their product is complete.
@@
It is proven in [bh] (II.9) that the Tits boundary $\partial_{T}(X_{1}\times X_{2})$ of the product of two complete CAT(0) spaces is the spherical join of the Tits boundaries of the factors. Therefore, it suffices to prove:
@@
The Tits boundary of $X_{p}(\mathbb{K})$ was studied by Duchesne [duc2013]: it is a thick spherical building, and all finite configurations of simplices are the same that appear in the spherical building of the finite-dimensional analogue $\mathrm{O}_{\mathbb{K}}(p,q)$ for sufficiently large $q$. Details on the structure of these buildings can be found in the book by Abramenko and Brown [buildings] (6.7). We show the following:
@@
We refer to [buildings] for details on the construction of the building. Here we only describe the structure of the fundamental apartment, which allows us to see that every simplex has diameter $<\pi/2$.
[…]
We now describe the structure of the simplices in the sphere, which is forced by symmetry. We can choose coordinates such that the sphere is the unit sphere in $\mathbb{R}^{p}$, for $1\leq i\leq p$ the subspace $\mathrm{Span}\{e_{i}\}$ is associated with the point with $i$-th coordinate equal to $1$, and the subspace $\mathrm{Span}\{f_{i}\}$ is associated with the point with $i$-th coordinate equal to $-1$. These points determine a subdivision of the sphere into $2^{p}$ super-simplices: each of them is the intersection of the sphere with the subset of $\mathbb{R}^{p}$ resulting from a specific choice of the signs of the $p$ coordinates. Each super-simplex corresponds to a maximal isotropic subspace, and is then subdivided barycentrically into $p!$ individual simplices. Each subset $A\subset\{e_{1},\ldots,e_{p},f_{1},\ldots,f_{p}\}$ with $\mathrm{Span}(A)$ totally isotropic is associated with the barycenter of the set of points associated with elements of $A$.
Clearly, all super-simplices are isometric, and all individual simplices are isometric. Let $\mathcal{S}$ be a super-simplex, without loss of generality,
@@
We assume without loss of generality $\sigma=id$ and show that the simplex $s_{id}$ has diameter $<\pi/2$. Let $x\in S^{p-1}$. The set of points in $S^{p-1}$ at distance $\geq\pi/2$ from $x$ can be defined by $S^{p-1}\cap\{y\in\mathbb{R}^{p}:\langle x,y\rangle\leq 0\}$, where we use the standard scalar product in $\mathbb{R}^{p}$. If we assume that there exist $x,y\in s_{id}$ with $d(x,y)\geq\pi/2$, they must satisfy $\langle x,y\rangle\leq 0$. As all their coordinates are non-negative and their first coordinate is at least $\sqrt{1/p}$, we have $\langle x,y\rangle\geq 1/p$, a contradiction. Hence, the diameter of $s_{id}$, and similarly the diameter of any simplex in the sphere, is strictly smaller than $\pi/2$.
@@
Suppose by contradiction that $\Sigma$ is the spherical join $A_{1}*A_{2}$. Then $A_{1},A_{2}$ are non-empty subsets of $\Sigma$ with the following properties (deduced from the construction of the spherical join, see [bh], I.5.13):
@@
Moreover, since $\Sigma$ is a CAT(1) space ([bh], II.10A.4), $A_{1}$ and $A_{2}$ must also be CAT(1) spaces ([bh], II.3.15). In particular they are $\pi$-geodesic.

### v2

@@ 3.2 Property (FH) and soluble normal subgroups
The main tool to prove Theorem 3.4 will be a form of Levi decomposition for parabolic subgroups of $H=\mathrm{Isom}(X)$, where a parabolic subgroup will be expressed as a semidirect product of a soluble normal subgroup and a Levi factor of lower rank. In this subsection, we prove a lemma that will allow us to restrict the action to the Levi factor, and thus to reduce the rank of the target space.
@@
Let $N^{\prime}:=[N,N]$ be the commutator subgroup of $N$, which has solubility length n1n-1, let $A=N/N^{\prime}$ be the abelianization of $N$, and let $\mathrm{ab}:N\rightarrow A$ be the natural projection (with kernel $N^{\prime}$). We recall that $Q$ acts on $N$ by the map $\phi:Q\rightarrow\mathrm{Aut}(N)$ that appears in the semidirect product and that associates to every $q\in Q$ the conjugation by $q$. Since the commutator subgroup $N^{\prime}$ is characteristic, it is preserved by the action of $Q$, which means that the action descends to the quotient $A$. We call $\phi_{\mathrm{ab}}$ this induced action. By hypothesis, $A$ is isomorphic (as a topological group) to the additive group of a Hilbert space, and the action $\phi_{\mathrm{ab}}$ of $Q$ preserves a scalar product in this Hilbert space.
@@
Using Property (FH), we find that the action of $\Gamma$ on $A$ fixes a point $a_{0}$, which means that if we conjugate $f$ by an element $n_{0}\in N$ with $\mathrm{ab}(n_{0})=a_{0}$ we get an action that fixes the neutral element of $A$, which means that a conjugate of the homomorphism $f$ has image contained in $N^{\prime}\rtimes Q$. Now, since $N^{\prime}$ has solubility length n1n-1, by inductive hypothesis this conjugate has image contained in a conjugate of $Q$, so in the end we find that $f$ has image contained in a conjugate of $Q$.
@@ 3.3 The induction in the finite-dimensional case
As the group $\Gamma$ has Property (F), we know that every action of $\Gamma$ on $X$ has to fix a point in $X$ or in $\partial X$. The first case is the thesis of our theorem, so from now on we will assume that we are in the second case, i. e. $\Gamma$ fixes $x\in\partial X$. This means that there is a continuous homomorphism $f:\Gamma\rightarrow H_{x}$, where $H_{x}<H=\mathrm{Isom}(X)$ is the parabolic subgroup defined by $H_{x}=\{h\in H:hx=x\}$.
In this subsection, we focus on the case where $X$ is finite-dimensional. We will use the classical theory of Lie groups and symmetric spaces of non-compact type, for which a good reference is the book by Eberlein [20]. We use results from (2.17) of that book, which we report here using our notation.
@@
Eberlein then defines $\mathfrak{n}_{x}$ as the Lie algebra of $N_{x}$ and shows that it is nilpotent (2.17.14), which means that the group $N_{x}$ is nilpotent in the algebraic sense ([28], Theorem 11.2.5). Moreover, as shown in the proof of the Iwasawa decomposition ([28], Theorem 13.3.8), $N_{x}$ is simply connected.
Now, if we have the homomorphism $f:\Gamma\rightarrow H_{x}$, we can compose it with the projection $\pi:H_{x}\rightarrow Z_{x}$ and get a homomorphism $\overline{f}:\Gamma\rightarrow Z_{x}$. In the proof of ([20], 2.17.5) it is shown that if $\gamma_{px}$ is a geodesic that passes through $p$ and points to $x$, the set $F(\gamma_{px})$ defined as the union of all geodesics parallel to $\gamma_{px}$ is preserved by the action of $Z_{x}$. Therefore, we have an action of $\Gamma$ on $F(\gamma_{px})$. It is shown in [20] (2.11.4) that $F(\gamma_{px})$ is a complete totally geodesic submanifold of $X$ and
@@
As $\Gamma$ acts on $F(\gamma_{px})$, the result of De Rham [40] implies that the action splits as the product of an action on $\mathbb{R}^{r}$ and an action on $F_{S}(\gamma_{px})$. Due to Property (FH), the action on $\mathbb{R}^{r}$ fixes a point; hence there exists an embedded copy of $F_{S}(\gamma_{px})$ that is preserved by $\Gamma$ and on which $\Gamma$ acts.
As $\Gamma$ has Property (F) it must fix a point in $F_{S}(\gamma_{px})$ or in its boundary. Since $F(\gamma_{px})$ is embedded in $X$ and since $r\geq 1$, the rank of $F_{S}(\gamma_{px})$ must be strictly lower than the rank of $X$, so we can use our inductive hypothesis and find that $\Gamma$ has to fix a point inside $F_{S}(\gamma_{px})$.
@@
To verify the hypotheses of Lemma 3.6, we need to see that each Abelian quotient that appears in the derived series of the soluble group $N_{x}$ is isomorphic to the additive group of a Hilbert space where the action of $Q$ preserves a scalar product. We know from [28] (Proposition 11.2.4) that the groups in the derived series of a connected Lie group $\mathcal{G}$ are integral subgroups (that is, they are generated by a subalgebra of the Lie algebra of $\mathcal{G}$), and that integral subgroups $\mathcal{H}$ of a simply connected soluble Lie group $\mathcal{G}$ are closed and simply connected, with $\mathcal{G}/\mathcal{H}$ isomorphic to $\mathbb{R}^{\dim{\mathcal{G}/\mathcal{H}}}$ ([28], Proposition 11.2.15). We can apply this to the Lie group $\mathcal{G}=N_{x}$, finding that the groups in the derived series are closed and simply connected, and each quotient is isomorphic to $\mathbb{R}^{n}$ for some $n$. We also know that the groups in the derived series are characteristic subgroups of $N_{x}$ (meaning that they are preserved by any automorphism of $N_{x}$), and hence the action of $Q$ given by the semidirect product $N_{x}\rtimes Q$ descends to an action on each quotient. Furthermore, since $Q$ is compact, each quotient has a scalar product that is preserved by this action of $Q$: it suffices to start from any scalar product and integrate it over $Q$ to obtain a $Q$-invariant one. This completes the proof.
@@ 3.4 The infinite-dimensional case
We are left with showing the inductive step for the infinite-dimensional case, where we have $X=X_{p}(\mathbb{K})$ for some $p\geq 1$ and $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$. Since $\Gamma$ acts continuously by isometries on $X$, it has a continuous homomorphism to the group $H=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, where $\mathrm{PO}_{\mathbb{R}}(p,\infty)=\mathrm{PO}(p,\infty)$, $\mathrm{PO}_{\mathbb{C}}(p,\infty)=\mathrm{PU}(p,\infty)$, and $\mathrm{PO}_{\mathbb{H}}(p,\infty)=\mathrm{PSp}(p,\infty)$. In fact, it is shown in [19] (Theorem 3.3) that $\mathrm{Isom}(X_{p}(\mathbb{K}))=\mathrm{PO}_{\mathbb{K}}(p,\infty)$, except in the case $\mathbb{K}=\mathbb{C}$ where $\mathrm{PU}(p,\infty)$ is a subgroup of index two in $\mathrm{Isom}(X_{p}(\mathbb{C}))$. In that case, we have a subgroup $\Gamma^{\prime}<\Gamma$ of index $\leq 2$ with a homomorphism to $\mathrm{PU}(p,\infty)$, and if we show that $\Gamma^{\prime}$ fixes a point in $X$, then this implies that $\Gamma$ has a finite orbit and hence fixes a point.
@@
Each point at infinity of $X_{p}(\mathbb{K})$ is associated with a flag of isotropic subspaces of the vector space $V$: this was shown in [16] (Proposition 6.1) in the real case, but is true in all cases, as observed in [19] (as a consequence of Remark 3.1 in [19]).
We consider the flag associated with the point at infinity $x$ (which is fixed by the action of $\Gamma$). Let $q\leq p$ be the dimension of the maximal subspace in the flag. We may choose the basis in such a way that this maximal subspace is exactly $\mathrm{Span}\{e_{1},\ldots,e_{q}\}$. Then we can change the order of the elements $(e_{q+1},e_{q+2},\ldots)$ of the basis in such a way that the sesquilinear form is now expressed by the matrix
@@
Now we consider the matrix associated with a transformation in $H_{x}$ with respect to this new basis and find some restrictions on its structure. We see it as a block matrix similar to the one expressing the sesquilinear form, i. e. with two blocks of dimension $q$ and one block of infinite dimension, both horizontally and vertically. This means that we are splitting the vector space $V$ into a direct sum $V_{1}\oplus V_{2}\oplus V_{3}$ where $V_{1}$ and $V_{2}$ have dimension $q$ and $V_{3}$ has infinite dimension, and each block of the matrix corresponds to a continuous linear map from $V_{i}$ to $V_{j}$ for some $i,j\in\{1,2,3\}$.
@@
In order to preserve the sesquilinear form, we must have
@@
The map sending a matrix as above to the block $B$ determines a homomorphism $N\rightarrow A$ where $A$ is the additive group of $q\times\infty$ matrices (i. e. linear maps from the $q$-dimensional vector space $V_{2}$ to the infinite-dimensional Hilbert space $V_{3}$), which is Abelian and isomorphic with the direct sum of $q$ Hilbert spaces of countable dimension, which is isomorphic with a single Hilbert space of countable dimension. The kernel of the map is the set of matrices in $N$ with $B=0$, which is isomorphic with the additive group of $q\times q$ matrices $Y$ that satisfy $Y+Y^{T}=0$, that is, the additive group of skew-symmetric matrices, isomorphic with $\mathbb{K}^{q(q-1)/2}$.
@@
What we need to show for our theorem is that any continuous homomorphism $f:\Gamma\rightarrow H_{x}$ has image contained in the stabilizer of a point in the symmetric space $X$. Consider the homomorphisms $f_{S}:\Gamma\rightarrow\mathrm{P}S$ and $f_{O}:\Gamma\rightarrow\mathrm{PO}_{\mathbb{K}}(p-q,\infty)$ defined by $f_{S}=\pi_{S}\circ f$ and $f_{O}=\pi_{O}\circ f$, where $\pi_{S}$ and $\pi_{O}$ are the natural projections of $H_{x}$ onto $\mathrm{P}S$ and $\mathrm{PO}_{\mathbb{K}}(p-q,\infty)$, respectively.
The homomorphism $f_{S}$ has image in $\mathrm{P}S<\mathrm{PGL}_{q}(\mathbb{K})$. As $\mathrm{PGL}_{q}(\mathbb{K})$ is a Lie group of rank q1<pq-1<p, we can use the inductive hypothesis of Theorem 3.4 and see that the image of $f_{S}$ fixes a point in the associated symmetric space. Up to conjugating the homomorphism (which is the same as changing the basis of $V_{1}$ and $V_{2}$ that we consider in our matrix decomposition), we can assume that the image of $f_{S}$ is contained in $\mathrm{PO}_{\mathbb{K}}(q)$.
@@ 3.5 Conclusion
We are now ready to prove Theorem 1.6, which implies Theorem 1.1 and Theorem 1.2.
Let $G=\mathrm{SL}_{3}(\mathbb{F})$ or $G=\mathrm{Sp}_{4}(\mathbb{F})$, with $\mathbb{F}$ a non-archimedean local field; in the case $G=\mathrm{Sp}_{4}(\mathbb{F})$ we assume that $\mathbb{F}$ has characteristic zero and its residue field $k$ has at least three elements. Let $\Gamma$ be a torsion-free cocompact lattice of $G$, which exists by ([6], Theorem A, Theorem 3.3) and ([24], Theorem 2.7). By Theorem 2.1, $\Gamma$ has Property (F).
The group $G$ has Kazhdan’s Property (T) ([4], 1.4, 1.5), and since this property is inherited by lattices ([4], Theorem 1.7.1), the lattice $\Gamma$ also has Kazhdan’s Property (T). By the Delorme-Guichardet Theorem ([4], Theorem 2.12.4), this means that these groups have Property (FH).
By Theorem 3.4, any continuous action by isometries of $\Gamma$ on a space $X\in\mathcal{X}$ has a fixed point in $X$. This proves Theorem 1.6.
@@ Appendix A Appendix
In this appendix we prove that the spaces $X_{p}(\mathbb{K})$, defined at the beginning of Section 3.1, are irreducible metric spaces, meaning that they are not isometric to the product of two metric spaces not reduced to a point.
Set $X=X_{p}(\mathbb{K})$. We suppose by contradiction that $X=X_{1}\times X_{2}$ for some metric spaces $X_{1},X_{2}$ not reduced to a point. First, we note that $X_{1},X_{2}$ must be geodesic spaces, since $X$ is a geodesic space and geodesics in a product project to geodesics in factors ([9], I.5.3). Moreover, geodesics in $X_{1},X_{2}$ must extend indefinitely, since geodesics in $X$ extend indefinitely. This implies that the boundary at infinity of $X_{1}$ and $X_{2}$ is non-empty. We also note that $X_{1},X_{2}$ must be CAT(0) spaces, as they are convex subsets of the CAT(0) space $X$, and that they must be complete, as their product is complete.
@@
It is proven in [9] (II.9) that the Tits boundary $\partial_{T}(X_{1}\times X_{2})$ of the product of two complete CAT(0) spaces is the spherical join of the Tits boundaries of the factors. Therefore, it suffices to prove:
@@
The Tits boundary of $X_{p}(\mathbb{K})$ was studied by Duchesne [16]: it is a thick spherical building, and all finite configurations of simplices are the same that appear in the spherical building of the finite-dimensional analogue $\mathrm{O}_{\mathbb{K}}(p,q)$ for sufficiently large $q$. Details on the structure of these buildings can be found in the book by Abramenko and Brown [1] (6.7). We show the following:
@@
We refer to [1] for details on the construction of the building. Here we only describe the structure of the fundamental apartment, which allows us to see that every simplex has diameter $\pi/2$.
The building is the flag complex of totally isotropic subspaces of a $(p+q)$-dimensional $\mathbb{K}$-vector space equipped with a sesquilinear form $\langle\cdot,\cdot\rangle$ of index $p$ (we assume $p<q$). The fundamental apartment is associated with a $(2p)$-tuple of linearly independent vectors $(e_{1},\ldots,e_{p},f_{1},\ldots,f_{p})$ such that $\langle e_{i},e_{j}\rangle=\langle f_{i},f_{j}\rangle=0$ for $1\leq i,j\leq p$, $\langle e_{i},f_{i}\rangle=1$ for $1\leq i\leq p$, and $\langle e_{i},f_{j}\rangle=0$ for $i\neq j$. The apartment is isometric to the sphere $S^{p-1}$ and is the flag complex of totally isotropic subspaces spanned by elements of the above $(2p)$-tuple.
Now we describe the structure of the simplices in the sphere, which is forced by symmetry. We can choose coordinates such that the sphere is the unit sphere in $\mathbb{R}^{p}$, for $1\leq i\leq p$ the subspace $\mathrm{Span}\{e_{i}\}$ is associated with the point with $i$-th coordinate equal to $1$, and the subspace $\mathrm{Span}\{f_{i}\}$ is associated with the point with $i$-th coordinate equal to 1-1. These points determine a subdivision of the sphere into $2^{p}$ super-simplices: each of them is the intersection of the sphere with the subset of $\mathbb{R}^{p}$ resulting from a specific choice of the signs of the $p$ coordinates. Each super-simplex corresponds to a maximal isotropic subspace and is then subdivided barycentrically into $p!$ individual simplices. Each subset $A\subset\{e_{1},\ldots,e_{p},f_{1},\ldots,f_{p}\}$ with $\mathrm{Span}(A)$ totally isotropic is associated with the barycenter of the set of points associated with elements of $A$.
Clearly, all super-simplices are isometric and all individual simplices are isometric. Let $\mathcal{S}$ be a super-simplex, without loss of generality,
@@
We assume without loss of generality $\sigma=id$ and show that the simplex $s_{id}$ has diameter $\pi/2$. Let $x\in S^{p-1}$. The set of points in $S^{p-1}$ at distance $\geq\pi/2$ from $x$ can be defined by $S^{p-1}\cap\{y\in\mathbb{R}^{p}:\langle x,y\rangle\leq 0\}$, where we use the standard scalar product in $\mathbb{R}^{p}$. If we assume that there exist $x,y\in s_{id}$ with $d(x,y)\geq\pi/2$, they must satisfy $\langle x,y\rangle\leq 0$. As all their coordinates are non-negative and their first coordinate is at least $\sqrt{1/p}$, we have $\langle x,y\rangle\geq 1/p$, a contradiction. Hence, the diameter of $s_{id}$, and similarly the diameter of any simplex in the sphere, is strictly smaller than $\pi/2$.
@@
Suppose by contradiction that $\Sigma$ is the spherical join $A_{1}*A_{2}$. Then $A_{1},A_{2}$ are non-empty subsets of $\Sigma$ with the following properties (deduced from the construction of the spherical join, see [9], I.5.13):
@@
Moreover, since $\Sigma$ is a CAT(1) space ([9], II.10A.4), $A_{1}$ and $A_{2}$ must also be CAT(1) spaces ([9], II.3.15). In particular they are $\pi$-geodesic.